If the side of a square increases by 10 find the increase in area will increase how much?​

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Answer :

Let the side of square= a

Area of square= s^2

given that side of square is increased by 10%.

area of new square = *1.1s) ^2

=1.21^2

total increase area= new square area-based of square

= 1.21s^2-s^2

0.21s^2

now percentage of increased area

=*0.21s^2/s^2)*100

=0.21/100

=21%

hence area of square will increase by 21%.


Thanks